OpenStudy (anonymous):

The selling price of x number of a certain stereo can be modeled by the function R(x) = 160x. The total cost of making x stereos is C(x) = 71x – 0.02x2. What is the percent markup for 31 stereos?

OpenStudy (ashleyisakitty):

R(31)=160*31=4960 C(31)= 71*31-0.02(31)^2=2181.78 (4960-2181.78)/2181.7=1.2734=127.3%

OpenStudy (anonymous):

@Ashleyisakitty Thank you!